frickeryz:

Multiply the following polynomials and simplify y(5y − 4) (9 + y)2

jhonyy9:

the last number of 2 not wan be an exponent ?

Imagine:

1 attachment
Imagine:

Let's try to use PEMDAS, Correct?

darkknight:

ASsuming it is an exponent @jhonyy9, alr so just distribute basically (9+y)^2 = y^2+18y+81 y(5y-4) = 5y^2-4y Now multiply (5y^2-4y) with (Y^2+18y+81) @imagine its multiplying the polynomials not pemdas but yes that always applies but the main thing is to "Multiply the following polynomials and simplify" if you read the question. And last thing simplify when u can so for example if you have x^2 and 3x^2 you can combine into 4x^2. Simplify all the like terms